The equation of tangents at point (0,0) to circle ,making intercepts of length 2a and 2b units on coordinate axis is(are)----------------
1
Expert's answer
2012-06-15T11:00:14-0400
(2a+0)^2+(2b+0)^2=1
4a^2+4b^2=1 or
a^2+b^2=1/4=0.25 - equation of a circle with radius 0.5